OpenStudy (anonymous):

On a map 1 inch = 5 miles. If a road is 130 miles long, how long will the road on the map be? inches

OpenStudy (anonymous):

I will give medal!

OpenStudy (vijeya3):

This is a logic based question.... see if i were to say that the cost of 5 pens id $20.What is the cost of 1 pen?Can u tell that?

OpenStudy (vijeya3):

@yasmeenK12 ?

OpenStudy (anonymous):

I was afk

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

0.25

OpenStudy (vijeya3):

oh?and how did you find that?may i know the method please?

OpenStudy (anonymous):

is it right?

OpenStudy (vijeya3):

nopes....see 5 pens=$20..then for 1 pen...won't u divide them?

OpenStudy (anonymous):

Oh its 4?

OpenStudy (vijeya3):

Yes !exactly...so did u understand the logic behind the pen example?

OpenStudy (anonymous):

yes

OpenStudy (vijeya3):

similarly,over here....u know that 1 inch=5miles...let the distance in inches be x....130 by 5...u'll get it in inches...can u try it now?

jhonyy9 (jhonyy9):

use the rule of 3 simple 1 inch --- 5 miles x inch ---- 130 miles ------------------- x = 130*1/5 = ? x will be the length of the road on the map

jhonyy9 (jhonyy9):

hope this will help you

OpenStudy (anonymous):

26?

OpenStudy (vijeya3):

yep!you got it girl!:D

OpenStudy (anonymous):

Thanks!

OpenStudy (vijeya3):

Never mind:) u can tag me whenever u need help..:D

OpenStudy (anonymous):

Can you help me with one more question?

OpenStudy (vijeya3):

sure:)

OpenStudy (anonymous):

Sophie deposited $500 and earned simple annual interest of 8%. How much money did Sophie have in her account after 2 years? $

OpenStudy (vijeya3):

do you know the formula for finding the simple interest?

OpenStudy (anonymous):

no

OpenStudy (vijeya3):

umm..okay...u knw what simple interest is?

OpenStudy (anonymous):

yes

OpenStudy (vijeya3):

okay...the formula for finding the simple interest is... SI=P*R*T/100 where P is the principal that is the amount u are depositing... R is the rate which here is 8% and T is the time for which u are depositing the money.. Can u find the SI now?

OpenStudy (vijeya3):

@yasmeenK12 ?

OpenStudy (anonymous):

I got the answer 580

OpenStudy (anonymous):

I don't need your help any more thank you!

OpenStudy (vijeya3):

Umm..okay:)